OpenStudy (anonymous):

Is there any way to do this integral by hand? My book says "use your computer" and gives an estimate, but we did some in class without a problem... (it's for surfaces of revolution): (The format is fixed in the first reply) integrate from 0 to pi/4: 2pi (tan x) \sqrt [1 + (sec^2 x)^2]

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/4} 2\pi (tanx)\sqrt{1+(\sec^2x)^2} dx\] There, that's a much better format

OpenStudy (precal):

yes but this one looks a little complicated, that is why the instructions stated to do it using the computer

OpenStudy (precal):

you could probably input it at www.wolframalpha.com

OpenStudy (precal):

MIT is known for their "Integration Bees". Also, in mathematics we have many books with tables of integral formulas.......Integrals can get very complicated really fast

OpenStudy (anonymous):

That's good to know :) I was afraid I was just totally missing some obvious step... I tried it at Wolfram|Alpha and the computation time expired :\

OpenStudy (precal):

I am not very good at inputting integration in wolf. You might want to repost the question but ask for help to input it in wolframalpha.